# Trap Identification in Molecular Charge Transport Networks

### v1

@@ Abstract
This paper introduces a method to identify traps in molecular charge transport networks as obtained by multiscale modeling of organic semiconductors. Depending on the materials, traps can be defect-like single molecules or clusters of several neighboring ones, and can have a significant impact on the dynamics of charge carriers. Our proposed method builds on the random walk model of charge dynamics on a directed, weighted graph, the molecular transport network. It comprises an effective heuristic to determine the number of traps or trap clusters based on the eigenvalues and eigenvectors of the random walk Laplacian matrix and uses subsequent spectral clustering techniques to identify these traps. In contrast to currently available methods, ours enables identification of trap molecules in organic semiconductors without having to explicitly simulate the charge dynamics. As a prototypical system we study an amorphous morphology of bathocuproine, a material with known high energetic disorder and charge trapping. Based on a first-principle multiscale model, we first obtain a reference charge transport network and then modify its properties to represent different trap characteristics. In contrast to currently available methods, our approach successfully identifies both single trap, multiple distributed traps, and a combination of a single-molecule trap and trap regions on an equal footing.
keywords:
First-principle Multiscale Model , Organic Semiconductors , Spectral Clustering , Charge Transport , Graph Theory
journal: Journal of Computational Physics
[…]
[1]organization=Department of Mathematics and Computer Science & Institute for Complex Molecular Systems, Eindhoven University of Technology, addressline=PO Box 513, city=Eindhoven, postcode=5600 MB, state=, country=Netherlands
1 Introduction
@@
A wide range of OSC applications [9, 10, 11, 12, 13] revealed that by tuning the number of charge carriers one can achieve controllable charge mobility. In trap dominated materials where the carrier number is greater than that of traps, only a portion of the carriers is captured by the traps and the remaining carriers can experience fast transport. For example, [14, 15] show that in the Gaussian disorder models used for the theoretical study of charge transport in OSC, when the carrier number is increased by two times, the mobility can increase by approximately 100 times.
Zooming into the molecular resolution, charge transport in OSCs is a sequence of transition events between the localized states [16, 17] and is modeled as a continuous time random walk (CTRW) process [18]. The transition rates of carriers depends on all the individual molecules’ geometries and relative orientation, which affect electronic structure properties such as the energy levels, electronic coupling elements between the molecules, and reorganization energies. Those quantities can be calculated from a first-principle multiscale model detailed in Section 2.
@@
At present, there are no methods for the identification of traps in molecular charge transport networks that perform reliably for all different trap types. Few attempts have been reported in identifying trap regions, or clusters, based on analyzing the actual simulated dynamics, e.g., via kinetic Monte Carlo (KMC) [20]. Qualitatively, once entered into such a trap region, the random walk (representing the charge dynamic of a single carrier) transitions mostly within it and escaping it is a rare event, making such KMC simulations very time-consuming. Two methods to accelerate KMC simulations which indirectly involve trap identification have previously been discussed. One is based on the (stochastic) watershed algorithm filling regions (”basins”) in the spatially resolved energy distribution [21]. This purely energy-based criterion does, however, not consider additional details of the factors influencing the molecular charge transport network. The second method [22] is based on a graph-theoretic decomposition (GD) and makes use of the fact that in the presence of trapping regions the Markov chain on the molecular charge transport network is nearly completely decomposable [23], allowing the associated graph to be partitioned into subgraphs. While this method takes the full information of the hopping-type dynamics into account, it is sensitive to the choice of parameters (related to, e.g., graph connectivity properties or transition rate ratios) and is not successful in identifying single trap nodes in the graph (as we will also discuss in Section 4.1.1).
@@
In this paper, we propose a new method that builds upon the idea of graph partitioning by using spectral clustering based on a specific type of Laplacian matrix of the graph [24, 25]. The aim of this method is to separate the graph into partitions, by minimizing a normalized cut cost function, such that the random walk processes rarely transitions between different partitions. While obtaining a minimized normalized cut is a NP-hard problem, a relaxed solution of this discrete optimization problem can be obtained from the eigenvectors of the random walk Laplacian matrix which will be introduced in Section 3. Our proposed method includes an effective heuristic for the determination of the number of traps or trap clusters based on these eigenvalues and eigenvectors of this random walk Laplacian and subsequent performing spectral clustering (using K-means clustering) to identify the traps. The former depends on a single threshold parameter for which we find an optimal choice neatly independent of the specific system. Using the charge transport network resulting from a multiscale model of an amorphous morphology of bathocuproine (BCP) [26], a molecular material with known high energetic disorder [27] and complex charge trapping behavior, we demonstrate that our approach successfully identifies both a single trap, multiple distributed traps, and a combination of a single-molecule trap and trap regions on an equal footing. We also find a strong relation between the cost function associated with the normalized cut and the charge-carrier dynamics simulated in a time-of-flight setup [28, 29], as well as the physical characteristics of the trap (regions).
@@
In what follows, Section 2 will introduce the elements of the first-principle multiscale model used to obtain the molecular charge transport network of BCP based on a combination of classical molecular dynamics (MD) with quantum electronic structure theory on the level of density-functional theory (DFT), and the calculation of the time-of-flight to assess charge-carrier dynamics from the model. In Section 3 we give the details of the spectral-clustering based trap identification method we propose in this work, including the determination of the cluster number and K-means clustering. The results of the application of this method to the BCP system and its modifications to cover different trap types is presented and discussed in Section 4. A brief conclusion and discussion concludes the paper.
2 Multiscale Model
@@
To create the reference molecular charge transport network of BCP, we employ a multiscale model which connects morphology simulations based on classical MD with quantum-classical methods to evaluate the Marcus rate for electron transfer between two molecules $i$ and $j$:
@@ 2.1 Molecular Dynamics
2.2 Electronic Structure Calculations
@@
$\left(-\frac{1}{2}\nabla^{2}_{\vec{r}}+v_{\text{ext}}(\vec{r})+v_{\text{H}}[\rho](\vec{r})+v_{\text{XC}}[\rho](\vec{r})\right)\phi_{l}(\vec{r})=H^{\text{KS}}\phi_{l}(\vec{r})=\epsilon_{l}\phi_{l}(\vec{r}),$ (3)
@@
2.3 Calculation of Marcus Rates
@@
$R_{ij}=U_{i}^{\text{nC}}-U_{i}^{\text{nN}}+U_{j}^{\text{cN}}-U_{j}^{\text{cC}}.$ (5)
Specifically for BCP, we obtain $R_{ij}=0.50\,\mathrm{eV}$.
The site energy $E_{i}=E_{i}^{\text{c}}-E_{i}^{\text{n}}$ is the difference between the total energies of the system in which molecule $i$ is carrying a charge or not, corresponding to the ionization potential in case of hole transport and the negative of the electron affinity in case of electron transport. These total energies in turn consist of different contributions associated with different physical mechanism, i.e.,
@@
where $J^{0}_{ij}=\langle\phi^{\text{HOMO}}_{i}|\hat{H}^{\text{KS}}_{\text{D}}|\phi^{\text{HOMO}}_{j}\rangle$, $e_{i}=\langle\phi^{\text{HOMO}}_{i}|\hat{H}^{\text{KS}}_{\text{D}}|\phi^{\text{HOMO}}_{i}\rangle$, $e_{j}=\langle\phi^{\text{HOMO}}_{j}|\hat{H}^{\text{KS}}_{\text{D}}|\phi^{\text{HOMO}}_{j}\rangle$, and $S_{ij}=\langle\phi^{\text{HOMO}}_{i}|\phi^{\text{HOMO}}_{j}\rangle$ with bra-ket notation. As the Hamiltonian of the dimer is diagonal in its eigenbasis $\left\{\phi_{\text{D}}(\vec{r})\right\}$ with eigenvalues $\left\{\epsilon_{D}\right\}$, so $H^{\text{KS}}_{\text{D}}=\text{diag}(\epsilon_{D})$, $J^{0}_{ij}$ can be calculated as $J^{0}_{ij}=\mathbf{p}_{i}^{\text{T}}\text{diag}(\epsilon_{\text{D}})\mathbf{p}_{j}$, where the vector $\mathbf{p}_{i}=\langle\phi_{i}^{\text{HOMO}}|\phi^{\text{HOMO}}_{\text{D}}\rangle$ and $\mathbf{p}_{j}=\langle\phi_{j}^{\text{HOMO}}|\phi_{\text{D}}^{\text{HOMO}}\rangle$ are the projections of the monomer orbitals $\phi_{i}^{\text{HOMO}},\phi_{j}^{\text{HOMO}}$ onto the dimer orbitals $\phi^{\text{HOMO}}_{\text{D}}$. Similarly, $e_{i(j)}=\mathbf{p}_{i(j)}^{\text{T}}\mathbf{p}_{i(j)}$ and $S_{ij}=\mathbf{p}_{i}^{\text{T}}\mathbf{p}_{j}$. All of these operations are performed in the basis set representation of the Kohn–Sham wave functions (see Sec. 2.2) as implemented in VOTCA. The resulting distribution of coupling elements for BCP is shown in Figure 10(b) in A.
2.4 Time-of-flight Calculation
After computation of the Markus rate from the multiscale model and defining the graph $\mathbf{G}=(\mathbf{V},\mathbf{W})$, the charge dynamics can be modeled as a continuous-time random walk process on the graph. To model the Source-Sink conditions, some vertices are used as Source to represent the electrode where charge carriers are injected, and some vertices as Sink where charge carriers are detected and time-of-flight (ToF) is recorded. Due to Pauli exclusion, each node can be occupied by at most one charge carrier, so a system with $N$ molecules and $N_{c}\leq N$ charge carriers has a total of (${N\choose N_{c}}$ occupation situations. Each of these occupation situations is considered as a state $\mathbf{s}$. If all the carriers’ occupations are in a Source node, the state is called a Source state, and if at least one of the Sink nodes is occupied, the state is called a Sink state. The transition rates between the states can be obtained from the matrix $\mathbf{W}:\omega_{ij}$, since the connectivity of states is encoded in the connectivity of the nodes (molecules), as detailed in  [29]. In particular, the transition rate $\Omega_{\mathbf{s}\mathbf{s}^{\prime}}$ from state $\mathbf{s}$ to $\mathbf{s}^{\prime}$ is:
@@ 2.4 Time-of-flight Calculation
with $N_{\text{source}}$ the number of source states.
3 Spectral Clustering Method
@@
3.1 Random Walk Laplacian Matrix and Graph Partitioning
@@
The problem is that solving Eq. 11 is known to be a NP hard problem. However, a relaxation of it can be solved using the so-called random walk Laplacian matrix. For a weighted graph $\mathbf{G}=(\mathbf{V},\mathbf{W})$ define the out-degree matrix by $\mathbf{D}:=D_{i,i}=\sum_{j}\omega_{ij}$, and the Laplacian matrix by $\mathbf{L}=\mathbf{D}-\mathbf{W}$. The random walk Laplacian matrix of a graph $\mathbf{G}$ is the given by
@@
subject to the constraint that $\mathbf{T}^{\prime}\mathbf{T}=\mathbf{I}$. The solution $\mathbf{T}^{\ast}$ to Eq. 14 is formed by the first $k$ eigenvectors of random walk Laplacian matrix $\mathbf{L}_{\text{rw}}$.
It should be noted that solving 14 does not yield a partition of the graph, as this is a relaxation of the true discrete optimization problem we wish to solve. To construct a partition in practice a K-means algorithm, introduced in the next subsection, is performed on the rows of solution $\mathbf{T}^{\ast}$ to 14.
3.2 K-means Clustering Algorithm
@@
6:Return cluster assignment $\mathcal{C}_{1},\cdots,\mathcal{C}_{k}$.
@@
3.3 Determination of the Cluster Number
@@
Starting from spectral clustering theory, we know that the multiplicity of the smallest eigenvalue 0 of a graph Laplacian equals the number of connected components in the graph $\mathbf{G}$. If $\mathbf{G}_{1},\cdots,\mathbf{G}_{k}$ are the connected components, then the eigenspace corresponding to eigenvalue 0 is spanned by the vectors $\{\mathbf{1}_{\mathbf{G}_{i}}\}_{i=1}^{k}$. Here $\mathbf{1}_{\mathbf{G}_{i}}$ is the vector where elements corresponding to the nodes in cluster $\mathbf{G}_{i}$ are one and the rest are 0. This shows that both the eigenvalues as well as the corresponding eigenvectors contain relevant information about the possible number of cluster.
Of course, in many applications, the graph is a single connected graph and the interesting clusters do not reside in disconnected components. Hence, we cannot simply use the multiplicity of the first eigenvector. Instead, a variety of ways to determine the number of clusters have been discussed previously. For instance, Fraley and Raftery proposed finding the number of clusters based on the log-likelihood of the data [50], while Still and Bialek suggested information-theoretic criteria based on the ratio of within cluster and between cluster similarity [51]. Tibshirani el at. used gap statistics on general data points to find the cluster number [52]. Specifically, in the heuristic eigengap method based on perturbation theory $k$ is chosen such that $\lambda_{1},\cdots,\lambda_{k}$ are very small and $\lambda_{k+1}$ is relatively large [25]. In practice, it can however be difficult to implement this heuristic as it is not well-defined what “relatively large“ is. This problem can clearly be seen in the eigenvalues of $\mathbf{L}_{\text{rw}}$ for the BCP model system, as shown on linear and log scale in Figure 1(a,b), respectively.
To overcome the non-obvious challenge of finding a good $k$, we notice that the ratio of the eigenvalue $\lambda_{i+1}/{\lambda_{i}}$ is large for the small indices $i$, as shown in Fig. 1(c). As the index $i$ becomes large, ${\lambda_{i+1}}/{\lambda_{i}}$ becomes small. This prompts us to utilize the index $\ell$ yielding the maximum ${\lambda_{\ell+1}}/{\lambda_{\ell}}$ as a good first guess for the number of clusters.
@@
All together, our method first chooses the index $\ell$ such that ${\lambda_{\ell+1}}/{\lambda_{\ell}}$ is the largest. Then using the eigenvectors corresponding to the first $\ell$ eigenvalues, we consider the elements sufficiently distinct from zero. This is controlled via a parameter $a$. For each of the first $\ell$ eigenvectors, we will look at the induced subgraph on those nodes whose entries are larger than $a$ in absolute value. We then collect the disconnected components of this induced subgraph as separate graphs. After this we have a collection of small subgraphs from which we will remove all subgraphs that are themselves a subgraph of another graph in this collection. This procedure leaves us with a list of disjoint subgraphs given by the first $\ell$ eigenvectors and we take $k$ to be the number of these graphs plus 1. The full procedure to determine the cluster number $k$ is summarized in Algorithm 2.
@@
5:Calculate $\ell=\text{argmax}(\frac{\lambda_{i+1}}{\lambda_{i}})$
6:Denote empty set $B=\emptyset$
7:for  $i=1,2,\cdots,\ell$  do
8:     Set node list $Q=\emptyset$
9:     for  $j=1,2,\cdots,N$  do
10:         if $\mathbf{u}_{ij}>a$ then
11:              add $j$ to $Q$
12:     Let $H_{1},\dots,H_{M}$ denote the disconnected components of the induced subgraph in $G$ on the nodes in $Q$.
13:     Update $B=B\cup\{H_{1},\dots,H_{M}\}$
14:for $H,H^{\prime}\in B$ do
15:     if $H\subseteq H^{\prime}$ then
16:         remove $H$ from $B$
17:$k=|B|+1$
@@
4 Results
We begin with a short semi-quantitive analysis of the effect of traps in a molecular charge transport network. As the name suggests, once a charge carrier encounters a trap, it will spend a significant amount of time in it and the observed charge dynamics will be slow (large ToF). However, if there are more carriers in the system than traps, one can expect that once all traps are filled, the remaining charge carriers are very mobile. Denote as $N_{t}$ the number of traps, with $N_{c}$ the number of charge carriers as in Section 2.4 and $\tau({N_{c}})$ the ToF depending on the number of charge carriers, evaluated, e.g., by Eqs. 9 and 10. Then one expects the ratio $\tau(N_{c})/\tau(N_{c}+1)$ to be large for $N_{c}=N_{t}$. Analysis of this ratio then provides a qualitative indication of the effective number of traps (depending on a definition of ”large”) in the transport network but not their location. In addition, the evaluation of $\tau(N_{c})$ for large number of carriers is computationally cumbersome [29].
To avoid such expensive calculations and still gain insight into which molecules correspond to the $N_{t}$ traps, one can consider all different subsets $Q(N_{t})$ of size $N_{t}$ of molecules in the network. We then consider the ToF of a single charge carrier $\tau(1)$ in the full system and in a system in which the subset $Q(N_{t})$ has been removed from the network, $\tau^{Q}(1)$. The removal of the nodes is motivated by the fact that in the case of very strong traps, carriers will not easily escape them and thus they will for $N_{c}>N_{t}$ be largely inaccessible to the mobile carriers. In this scenario, one can inspect the ratio $\tau(1)/\tau^{Q}(1)$ and identify traps from its ”large” values.

### v2

# Automatic Identification of Traps in Molecular Charge Transport Networks of Organic Semiconductors

@@ Abstract
This paper introduces a method to identify traps in molecular charge transport networks as obtained by multiscale modeling of organic semiconductors. Depending on the materials, traps can be defect-like single molecules or clusters of several neighboring ones, and can have a significant impact on the dynamics of charge carriers. Our proposed method builds on the random walk model of charge dynamics on a directed, weighted graph, the molecular transport network. It comprises an effective heuristic to determine the number of traps or trap clusters based on the eigenvalues and eigenvectors of the random walk Laplacian matrix and uses subsequent spectral clustering techniques to identify these traps. In contrast to currently available methods, ours enables identification of trap molecules in organic semiconductors without having to explicitly simulate the charge dynamics and is applicable to a variety of energy- or topology-based traps in homomolecular or mixed systems with or without detailed-balance. As a prototypical system we simulate an amorphous morphology of bathocuproine, a material with known high energetic disorder and charge trapping. Based on a first-principle multiscale model, we first obtain a reference charge transport network and then purposefully modify its properties to represent different trap characteristics. In contrast to currently available methods, our approach successfully identifies both single trap, multiple distributed traps, and a combination of a single-molecule trap and trap regions on an equal footing.
I Introduction
@@
A wide range of OSC applications [9, 10, 11, 12, 13] revealed that by tuning the number of charge carriers one can achieve controllable charge mobility. In trap dominated materials where the carrier number is greater than that of traps, only a portion of the carriers is captured by the traps and the remaining carriers can experience fast transport. For example, [14, 15] show that in the Gaussian disorder models used for the theoretical study of charge transport in OSC, when the carrier number is increased by a factor two, the mobility can increase by approximately 100 times. Intriguingly, traps also play an important role for the conduction mechanism in inorganic electronic materials, e.g., in the context of trap-assissted tunneling [16, 17].
Zooming into the molecular resolution, charge transport in OSCs is a sequence of transition events between the localized states [18, 19] and is modeled as a continuous time random walk (CTRW) process [20]. The transition rates of carriers depends on all the individual molecules’ geometries and relative orientation, which affect electronic structure properties such as the energy levels, electronic coupling elements between the molecules, and reorganization energies. Those quantities can be calculated from a first-principle multiscale model detailed in Section II.
@@
At present, there are no methods for the identification of traps in molecular charge transport networks that perform reliably for all different trap types. Few attempts have been reported in identifying trap regions, or clusters, based on analyzing the actual simulated dynamics, e.g., via kinetic Monte Carlo (KMC) [22]. Qualitatively, once entered into such a trap region, the random walk (representing the charge dynamic of a single carrier) transitions mostly within it and escaping it is a rare event, making such KMC simulations very time-consuming. Two methods to accelerate KMC simulations which indirectly involve trap identification have previously been discussed. One is based on the (stochastic) watershed algorithm filling regions (”basins”) in the spatially resolved energy distribution [23]. This purely energy-based criterion does, however, not consider additional details of the factors influencing the molecular charge transport network. The second method [24] is based on a graph-theoretic decomposition (GD) and makes use of the fact that in the presence of trapping regions the Markov chain on the molecular charge transport network is nearly completely decomposable [25], allowing the associated graph to be partitioned into subgraphs. While this method takes the full information of the hopping-type dynamics into account, it is sensitive to the choice of parameters (related to, e.g., graph connectivity properties or transition rate ratios) and is not successful in identifying single trap nodes in the graph (as we will also discuss in Section IV.1.1).
@@
In this paper, we propose a new method that builds upon the idea of graph partitioning by using spectral clustering based on a specific type of Laplacian matrix of the graph [26, 27]. The aim of this method is to separate the graph into partitions, by minimizing a normalized cut cost function, such that the random walk processes rarely transitions between different partitions. While obtaining a minimized normalized cut is a NP-hard problem, a relaxed solution of this discrete optimization problem can be obtained from the eigenvectors of the random walk Laplacian matrix which will be introduced in Section III. Our proposed method includes an effective heuristic to determine the number of traps or trap clusters based on these eigenvalues and eigenvectors of this random walk Laplacian and subsequent performing spectral clustering (using K-means clustering) to identify the traps. The former depends on a single threshold parameter for which we find an optimal choice nearly independent of the specific system.
Using the charge transport network resulting from a multiscale model of an amorphous morphology of bathocuproine (BCP) [28], a molecular material with known high energetic disorder [29], low charge mobility [30], and complex charge trapping behavior, as a baseline with subsequent modifications we demonstrate that our approach successfully identifies energy-based single traps, multiple distributed traps, and a combination of a single-molecule trap and trap regions on an equal footing. It is furthermore shown that the approach is robust under application of external electric fields and also detects topology-based traps and traps in mixed-molecular materials that are not energy based without modifications to the algorithm. We also find a strong relation between the cost function associated with the normalized cut and the charge-carrier dynamics simulated in a time-of-flight setup [31, 32], as well as the physical characteristics of the trap (regions).
@@
In what follows, Section II will introduce the elements of the first-principle multiscale model used to obtain the molecular charge transport network of BCP based on a combination of classical molecular dynamics (MD) with quantum electronic structure theory on the level of density-functional theory (DFT), and the calculation of the time-of-flight to assess charge-carrier dynamics of the model. In Section III we give the details of the spectral-clustering based trap identification method we propose in this work, including the determination of the cluster number and K-means clustering. The results of the application of this method to the BCP system and its modifications to cover different trap types is presented and discussed in Section IV. A brief conclusion and discussion concludes the paper.
II Multiscale Model
@@
To create the baseline molecular charge transport network of BCP, we employ a multiscale model which connects morphology simulations based on classical MD with quantum-classical methods to evaluate the Marcus rate for electron transfer between two molecules $i$ and $j$:
@@ II.1 Molecular Dynamics
II.2 Electronic Structure Calculations
@@
$\begin{split}\left(-\frac{1}{2}\nabla^{2}_{\vec{r}}+v_{\text{ext}}(\vec{r})+v_{\text{H}}[\rho](\vec{r})+v_{\text{XC}}[\rho](\vec{r})\right)\phi_{l}(\vec{r})\\ =H^{\text{KS}}\phi_{l}(\vec{r})=\epsilon_{l}\phi_{l}(\vec{r}),\end{split}$ (3)
@@
II.3 Calculation of Marcus Rates
@@
$R_{ij}=R_{i}^{\text{n}}+R_{j}^{\text{c}}=U_{i}^{\text{nC}}-U_{i}^{\text{nN}}+U_{j}^{\text{cN}}-U_{j}^{\text{cC}}.$ (5)
Specifically for BCP, we obtain the value $R_{ij}=0.49\,\mathrm{eV}$, which is considered a single-molecule quantity in homo-molecular materials.
For the remaining quantities that depend on the material morphology (site energies and electronic coupling), special care is taken to avoid discrepancies between the molecular structures obtained via classical MD and the DFT calculations. To remove bond length fluctuations introduced by molecular dynamics simulations, as they are already integrated out in the derivation of the rate expression, molecular fragments with rigid, planar $\pi$ systems are substituted by DFT-optimized versions [19]. The site energy $E_{i}=E_{i}^{\text{c}}-E_{i}^{\text{n}}$ is the difference between the total energies of the system in which molecule $i$ is carrying a charge or not, corresponding to the ionization potential in case of hole transport and the negative of the electron affinity in case of electron transport. These total energies in turn consist of different contributions associated with different physical mechanism, i.e.,
@@
where $J^{0}_{ij}=\langle\phi_{i}|\hat{H}^{\text{KS}}_{\text{D}}|\phi_{j}\rangle$, $e_{i}=\langle\phi_{i}|\hat{H}^{\text{KS}}_{\text{D}}|\phi_{i}\rangle$, $e_{j}=\langle\phi_{j}|\hat{H}^{\text{KS}}_{\text{D}}|\phi_{j}\rangle$, and $S_{ij}=\langle\phi_{i}|\phi_{j}\rangle$ with bra-ket notation. The Hamiltonian of the dimer, $H^{\text{KS}}_{\text{D}}$ (see Eq. 3), is diagonal in its eigenbasis $\left\{|\phi^{\text{D}}_{k}\rangle\right\}$ with eigenvalues $\left\{\epsilon^{\text{D}}_{k}\right\}$, so $H^{\text{KS}}_{\text{D}}=\text{diag}(\epsilon^{\text{D}})$. With the projections of the monomer functions on the dimer eigenbasis, i.e., $p_{ik}=\langle\phi_{i}|\phi^{\text{D}}_{k}\rangle$ and $p_{jk}=\langle\phi_{j}|\phi^{\text{D}}_{k}\rangle$, $J^{0}_{ij}$ can be calculated as $J^{0}_{ij}=\mathbf{p}_{i}^{\text{T}}\text{diag}(\epsilon^{\text{D}})\mathbf{p}_{j}$. Similarly, $e_{i(j)}=\mathbf{p}_{i(j)}^{\text{T}}\mathbf{p}_{i(j)}$ and $S_{ij}=\mathbf{p}_{i}^{\text{T}}\mathbf{p}_{j}$. All of these operations are performed in the basis set representation of the Kohn–Sham wave functions (see Sec. II.2) as implemented in VOTCA. The resulting distribution of coupling elements for BCP is shown in Figure 12(b) in A.
II.4 Time-of-flight Calculation
After computation of the Marcus rate from the multiscale model and defining the graph $\mathbf{G}=(\mathbf{V},\mathbf{W})$, the charge dynamics can be modeled as a continuous-time random walk process on the graph. To model the Source-Sink conditions, some vertices are used as Source to represent the electrode where charge carriers are injected, and some vertices as Sink where charge carriers are detected and time-of-flight (ToF) is recorded. Specifically, we consider as source vertices the molecules with a COM component $r_{i}^{x}<0.5\,\mathrm{nm}$, and as sink vertices those with $r_{i}^{x}>\max(r_{i}^{x})-0.5\,\mathrm{nm}$. As a result, the graph retains the cyclic boundary conditions only in the Cartesian $y$- and $z$-directions.
In the model, each node can be occupied by at most one charge carrier, so a system with $N$ molecules and $N_{c}\leq N$ charge carriers has a total of (${N\choose N_{c}}$ occupation situations. While spin is not explicitly taken into account, the assumption of such a Pauli-like exclusion is a computationally effective way to avoid the calculation of explicit long-range Coulomb repulsion between electrons at different sites. Note that the formation of double excitations or bi-polarons is not included in the main transport dynamics model and requires different rate formulations. Each of these occupation situations is considered as a state $\mathbf{s}$. If all the carriers’ occupations are in a Source node, the state is called a Source state, and if at least one of the Sink nodes is occupied, the state is called a Sink state. The transition rates between the states can be obtained from the matrix $\mathbf{W}:\omega_{ij}$, since the connectivity of states is encoded in the connectivity of the nodes (molecules), as detailed in  [32]. In particular, the transition rate $\Omega_{\mathbf{s}\mathbf{s}^{\prime}}$ from state $\mathbf{s}$ to $\mathbf{s}^{\prime}$ is:
@@ II.4 Time-of-flight Calculation
with $N_{\text{source}}$ the number of source states. Note that the above model does not include a constant supply of charge carriers at the source but instead corresponds to a simulation with a short, pulsed creation of charge carriers, equivalent to KMC or Master Equation based models [32].
III Spectral Clustering Method
@@
III.1 Random Walk Laplacian Matrix and Graph Partitioning
@@
The problem is that solving Eq. 11 is known to be a NP hard problem. However, a relaxation of it can be solved using the so-called random walk Laplacian matrix. For a weighted graph $\mathbf{G}=(\mathbf{V},\mathbf{W})$ define the out-degree matrix by $\mathbf{D}:=D_{i,i}=\sum_{j}\omega_{ij}$, and the Laplacian matrix by $\mathbf{L}=\mathbf{D}-\mathbf{W}$. The random walk Laplacian matrix of a graph $\mathbf{G}$ is given by
@@
subject to the constraint that $\mathbf{T}^{\prime}\mathbf{T}=\mathbf{I}$. The solution $\mathbf{T}^{\ast}$ to Eq. 14 is formed by the first $k$ eigenvectors of the random walk Laplacian matrix $\mathbf{L}_{\text{rw}}$.
It should be noted that solving 14 does not yield a partition of the graph, as this is a relaxation of the true discrete optimization problem we wish to solve. To construct a partition in practice a K-means algorithm, introduced in the next subsection, is performed on the rows of the solution $\mathbf{T}^{\ast}$ to 14.
III.2 K-means Clustering Algorithm
@@
6:end while
7:Return cluster assignment $\mathcal{C}_{1},\cdots,\mathcal{C}_{k}$.
@@
III.3 Determination of the Cluster Number
@@
Starting from spectral clustering theory, we know that the multiplicity of the smallest eigenvalue 0 of a graph Laplacian equals the number of connected components in the graph $\mathbf{G}$. If $\mathbf{G}_{1},\cdots,\mathbf{G}_{k}$ are the connected components, then the eigenspace corresponding to eigenvalue 0 is spanned by the vectors $\{\mathbf{1}_{\mathbf{G}_{i}}\}_{i=1}^{k}$. Here $\mathbf{1}_{\mathbf{G}_{i}}$ is the vector where elements corresponding to the nodes in cluster $\mathbf{G}_{i}$ are one and the rest are 0. This shows that both the eigenvalues as well as the corresponding eigenvectors contain relevant information about the possible number of clusters.
Of course, in many applications, the graph is a single connected graph and the interesting clusters do not reside in disconnected components. Hence, we cannot simply use the multiplicity of the first eigenvector. Instead, a variety of ways to determine the number of clusters have been discussed previously. For instance, Fraley and Raftery proposed finding the number of clusters based on the log-likelihood of the data [53], while Still and Bialek suggested information-theoretic criteria based on the ratio of within cluster and between cluster similarity [54]. Tibshirani el at. used gap statistics on general data points to find the cluster number [55]. Specifically, in the heuristic eigengap method based on perturbation theory $k$ is chosen such that $\lambda_{1},\cdots,\lambda_{k}$ are very small and $\lambda_{k+1}$ is relatively large [27]. In practice, it can however be difficult to implement this heuristic as it is not well-defined what “relatively large“ is. This problem can clearly be seen in the eigenvalues of $\mathbf{L}_{\text{rw}}$ for the BCP model system, as shown on linear and log scale in Figure 1(a) and Figure 1 (b), respectively.
To overcome the non-obvious challenge of finding a good $k$, we notice that the ratio of the eigenvalue $\lambda_{i+1}/{\lambda_{i}}$ is large for small indices $i$, as shown in Fig. 1(c). As the index $i$ becomes large, ${\lambda_{i+1}}/{\lambda_{i}}$ becomes small. This prompts us to utilize the index $\ell$ yielding the maximum ${\lambda_{\ell+1}}/{\lambda_{\ell}}$ as a good first guess for the number of clusters.
@@
All together, our method first chooses the index $\ell$ such that ${\lambda_{\ell+1}}/{\lambda_{\ell}}$ is the largest. Then using the eigenvectors corresponding to the first $\ell$ eigenvalues, we consider the elements sufficiently distinct from zero. This is controlled via a parameter $a$. For each of the first $\ell$ eigenvectors, we will look at the induced subgraph on those nodes whose entries are larger than $a$ in absolute value. We then collect the disconnected components of the corresponding induced subgraph as separate graphs. After this we have a collection of small subgraphs from which we will remove all subgraphs that are themselves a subgraph of another graph in this collection. This procedure leaves us with a list of disjoint subgraphs given by the first $\ell$ eigenvectors and we take $k$ to be the number of these graphs plus 1. The full procedure to determine the cluster number $k$ is summarized in Algorithm 2.
@@
5:end for
6:Calculate $\ell=\text{argmax}(\frac{\lambda_{i+1}}{\lambda_{i}})$
7:Denote empty set $B=\emptyset$
8:for  $i=1,2,\cdots,\ell$  do
9:     Set node list $Q=\emptyset$
10:     for  $j=1,2,\cdots,N$  do
11:         if $\mathbf{u}_{ij}>a$ then
12:              add $j$ to $Q$
13:         end if
14:     end for
15:     Let $H_{1},\dots,H_{M}$ denote the disconnected components of the induced subgraph in $G$ on the nodes in $Q$.
16:     Update $B=B\cup\{H_{1},\dots,H_{M}\}$
17:end for
18:for $H,H^{\prime}\in B$ do
19:     if $H\subseteq H^{\prime}$ then
20:         remove $H$ from $B$
21:     end if
22:end for
23:$k=|B|+1$
@@
IV Results
We begin with a short semi-quantitive analysis of the effect of traps in a molecular charge transport network. As the name suggests, once a charge carrier encounters a trap, it will spend a significant amount of time in it and the observed charge dynamics will be slow (large ToF). However, if there are more carriers in the system than traps, one can expect that once all traps are filled, the remaining charge carriers are very mobile. Denote as $N_{t}$ the number of traps, with $N_{c}$ the number of charge carriers as in Section II.4 and $\tau({N_{c}})$ the ToF depending on the number of charge carriers, evaluated, e.g., by Eqs. 9 and 10. Then one expects the ratio $\tau(N_{c})/\tau(N_{c}+1)$ to be large for $N_{c}=N_{t}$. Analysis of this ratio then provides a qualitative indication of the effective number of traps (depending on a definition of ”large”) in the transport network but not their location or their physical nature. In addition, the evaluation of $\tau(N_{c})$ for large number of carriers is computationally cumbersome [32].
To avoid such expensive calculations and still gain insight into which molecules correspond to the $N_{t}$ traps, one can consider all different subsets $Q(N_{t})$ of size $N_{t}$ of molecules in the network. We then consider the ToF of a single charge carrier $\tau(1)$ in the full system and in a system in which the subset $Q(N_{t})$ has been removed from the network, $\tau^{Q}(1)$. The removal of the nodes is motivated by the fact that in the case of very strong traps, carriers will not easily escape them and thus they will for $N_{c}>N_{t}$ be largely inaccessible to the mobile carriers. In this scenario, one can inspect the ratio $\tau(1)/\tau^{Q}(1)$ and identify traps from its ”large” values. Such change in ToF provides a qualitative and system-agnostic way to identify traps in complex molecular charge transport networks, where site energy or other static properties alone may not suffice.